OpenStudy (anonymous):

Jane bought a car for $30,303. The value of the car depreciated at a constant rate per year. The table below shows the value of the car after the first and second years: Year 1 2 Value (in dollars) 25,757.55 21,893.92 Which function best represents the value of the car after t years?

OpenStudy (anonymous):

f(t) = 25,757.55(0.85)t f(t) = 30,303(0.85)t f(t) = 30,303 (0.15)t f(t) = 25,757.55 (0.15

OpenStudy (anonymous):

f(t) = 25,757.55(0.85)t f(t) = 30,303(0.85)t f(t) = 30,303 (0.15)t f(t) = 25,757.55 (0.15

OpenStudy (godlovesme):

is that an option?

OpenStudy (anonymous):

the options yes

OpenStudy (godlovesme):

sorry idk this :( @Nnesha can u help

OpenStudy (anonymous):

@bruhhh

OpenStudy (anonymous):

@billj5

OpenStudy (godlovesme):

@SolomonZelman @Directrix

OpenStudy (anonymous):

@freckles

Nnesha (nnesha):

okay  now first amount was 30303 so solve your answer choices and see which answer choice give u 25,757.55 <--this number